3(x+2)(x-3)(x+9)=0
Answer by venugopalramana(3286)

3(x+2)(x-3)(x+9)=0
IF A PRODUCT IS ZERO THEN ONE OF THE FACTORS SHOULD BE ZERO.
HENCE WE GET
X+2=0...OR
X-3=0....OR
X+9=0...THAT IS
X=-2 OR 3 OR -9 ..SO THEY ARE THE POSSIBLE SOLUTIONS FOR THIS EQUATION.